OpenStudy (anonymous):

solve the following differential eqn: x'-3x+x^2=0, if x=1 when t=0, determine the value of x when t=1

OpenStudy (turingtest):

this one is separable

OpenStudy (anonymous):

its not non-homo? i mean if i take x^2 to RHS ... then solve LHS as a compl soln... then RHS as particular soln???

OpenStudy (turingtest):

perhaps you are right, sorry. one sec please...

OpenStudy (turingtest):

this is linear, and so can be with an integrating factor I believe

OpenStudy (turingtest):

no I am confused by the x I think :P hm... bernoulli I guess I'm not so good at those

OpenStudy (anonymous):

not good?

OpenStudy (turingtest):

I am not good at bernoulli problems, but others are

OpenStudy (sirm3d):

it is bernoulli's

OpenStudy (turingtest):

thought so. well I'll leave it to you then :)

OpenStudy (anonymous):

wats bernoulli?

OpenStudy (sirm3d):

the trick with bernoulli's equation \[x \prime + p(t) x=q(t)x^n\] is to divide everything by x^n

OpenStudy (anonymous):

wow thanks @TuringTest @sirm3d my x^n is my x^2?

OpenStudy (sirm3d):

yes, you're right

OpenStudy (anonymous):

x'/x^2 -3x/x^2=-1

OpenStudy (anonymous):

x'/x^2-3/x=-1

OpenStudy (sirm3d):

or \[x^{-2}x \prime -3x^{-1}=-1\] where you sub \[v=x^{-1} \] . this should get you going.

OpenStudy (anonymous):

i'm confused... v^(-1)x'-3v=-1?

OpenStudy (sirm3d):

whenever a sub is made in a DE, the derivative is generally expected of the sub. here \[v=x^{-1}\] so \[v \prime = -x^{-2} x \prime\]

OpenStudy (sirm3d):

can you do it from here?

OpenStudy (anonymous):

:(

OpenStudy (sirm3d):

let me show it then \[x^{-2}x \prime-3x^{-1}=-1\] by sub \[v=x^{-1}\] and \[v \prime = -x^{-2}x \prime\] we get \[-v \prime-3v=-1\]

OpenStudy (sirm3d):

@GeekChic_ are you still with me in this problem?

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

@sirm3d u ther?

OpenStudy (anonymous):

now that i got to v'-3v=-1 i solve it normally? complementary n particular solns?

OpenStudy (anonymous):

Yc=Ae^(3v) and Yp=1/2 gen soln Y=Ae^(3v)+1/2

OpenStudy (sirm3d):

wait, a kid is messing with my mouse.

OpenStudy (sirm3d):

you can use an integrating factor. the integrating factor for \[y\prime+p(x) y=q(x)\] is \[v=\int\limits_{}^{}p(x)dx\] and the solution to the DE is \[vy=\int\limits_{}^{}vq(x)dx\]

OpenStudy (sirm3d):

let me correct the integrating factor. it should be \[v=e^{\int\limits_{}^{}p(x)dx}\] do you still remember this method of solving DE, @GeekChic_ ?

OpenStudy (anonymous):

noooooo

OpenStudy (sirm3d):

oh no. one of the classes of 1st order linear DE is \[y\prime + p(x)y=q(x)\] and its solution is by using an integrating factor that will make the DE exact. the integrating factor turns out to be \[e^{\int\limits_{}^{}p(x)dx}\]

OpenStudy (sirm3d):

@GeekChic_ BUT you can use \[v_c =Ae^{-3x}\] and \[v_p = \frac{ 1 }{ 3 }\]

OpenStudy (sirm3d):

sorry it's not x but t as the independent variable. the general solution is \[v=Ae^{-3t} + \frac{ 1 }{ 3}\] or in terms of x, \[x^{-1}=Ae^{-3t}+\frac{1}{3}\] sub t=0, x=1 to find A, then find x when t=1.

OpenStudy (sirm3d):

just one thing. the particular solution Yp may be difficult to find for the general function q(x) on the DE \[y\prime + p(x)y = q(x)\], so it is advisable to learn the method of integrating factor posted above.

OpenStudy (anonymous):

wow... u are soooo good!!!! thank u sooo much

OpenStudy (sirm3d):

yw. =)

OpenStudy (anonymous):

hi how are you?

OpenStudy (anonymous):

@ sirmed, can do the beginning again plz

OpenStudy (sirm3d):

which part ^ is not clear to you?

OpenStudy (anonymous):

ok never mind lo0l.. i just did it now and its ok now lol..thnx anyway.. :D

OpenStudy (anonymous):

at first i did like this x'-3x= -x^2 will that also work?

OpenStudy (sirm3d):

that's the first step, all right.

OpenStudy (anonymous):

i didnt divide it to x^2 but instead directly work into sub x=u^-1, x'=-u^-2, will x^2=u^-2?

OpenStudy (sirm3d):

that's how bernoulli's equation is solved, the substitution \[u=x^{1-n}\]

OpenStudy (anonymous):

ok i know that but how about like this? x'-3x= -x^2, x=u^-1, x'=-u^-2, will x^2=u^-2? -u^-2 -3u^-1 = - u^-2 ?? will this work on the right side?

OpenStudy (anonymous):

-u^-2 u'-3u^-1 = - u^-2 ?? will this work on the right side?

OpenStudy (sirm3d):

same result. after all, if u=x^-1, then x=u^-1

OpenStudy (anonymous):

ok if continuing it would be -u^-2 u'-3u^-1 = - u^-2 ----------------------- -u^-2 u'-3u=1 ? and your solution is -u'-3u=-1 its the negative signs?

OpenStudy (sirm3d):

each term of your equation has a negative sign, so all terms are positive after division by -u^-2

OpenStudy (anonymous):

ooops its +3u..lol :D sorry i didnt see that hahaha

OpenStudy (anonymous):

ok then thank u very much ... lol :D

OpenStudy (sirm3d):

therefore we ended with the same equation. YW.

OpenStudy (anonymous):

sometimes for just a little bit thing, everything messede up..lol... :D thnx a lot and have fun solving now...